# Planckian bound on quantum dynamical entropy

### v3

@@ Abstract
We introduce a simple definition of dynamical entropy for quantum systems under continuous monitoring, inspired by Connes, Narnhofer and Thirring. It quantifies the amount of information gained about the initial condition. A nonzero entropy rate can be obtained by monitoring the thermal fluctuation of an extensive observable in a generic many-body system (away from classical or large $N$ limit). We explicitly compute the entropy rate in the thermodynamic and long-time limit, in terms of the two-point correlation functions. We conjecture a universal Planckian bound for the entropy rate. Related results on the purification rate are also obtained.
@@
Defining a dynamical entropy for quantum systems has been an outstanding question since KS [8]. A few different proposals have been put forward [9, 10, 11, 12, 13, 14, 15, 16, 17], and concrete calculations have been performed for simple systems such as free quantum gases [18, 14, 19]. However, it remains unknown whether quantum dynamical entropy can characterize chaos and complexity in many-body quantum systems (see [20] for a recent survey). This question is timely as many-body quantum chaos has been an intensively pursued topic recently [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. One interesting approach is the out-of-time order correlation (OTOC) [37], which generalizes the classical butterfly effect to quantum systems. Remarkably, its growth rate (Lyapunov exponent) obeys a Planckian bound at low temperatures. The Planckian bound showcases a universal limit on dynamical chaos due to quantum effects, and is famously saturated by strongly interacting holographic models [38, 39, 40, 41, 42]; see also [43, 44, 45, 46, 47, 48] for other “Planckian bounds”. However, OTOCs do not have a well-defined exponential growth in generic many-body quantum systems with local interactions, that is, away from semiclassical or large N limits. Meanwhile, quantum chaos diagnosis inspired by random matrix theory [49, 50, 51] probe time scales that diverge in the thermodynamic limit.
In this Letter, we propose a simplified version of Connes-Narnhofer-Thirring (CNT) quantum dynamical entropy [10]. We show that it has a well-defined linear growth, in generic quantum many-body systems at thermal equilibrium, and at $O(1)$ time in the thermodynamic limit. Moreover, we conjecture that the growth rate satisfies a Planckian bound: in units with $\hbar=k_{B}=1$,
@@
where $T$ is the temperature and $c$ is a universal constant. These results suggest the dynamical entropy as an interesting probe of low-temperature many-body quantum chaos.
In what follows, we first define the quantum dynamical entropy using a setup of purification by continuous monitoring [19, 52, 53], see eq. (Planckian bound on quantum dynamical entropy) and Fig. 1 for illustration.. Compared to the classical KS entropy, the CNT entropy accounts for the back-reaction of quantum measurements, which prevents future measurements from gaining information on the initial state. We then argue that, as a result, a positive dynamical entropy rate can be obtained only by monitoring the fluctuation of an extensive quantity, see (9)-(10) below. As a main technical result, we derive an exact formula (12) that computes the asymptotic dynamical entropy growth rate. We then argue for the Planckian bound using the exact rate formula. We finally present analogous results on purification.
@@
Setup. Consider a quantum system $A$, whose initial state is mixed, and given by the density matrix $\rho$. We shall purify $\rho$ as the reduced density matrix of the pure state $\sqrt{\rho}$ on $A\bar{A}$, where $\bar{A}$ is an identical copy of $A$. (The reduced density matrix of $\bar{A}$ is also $\rho$.) We shall by default be concerned with the case where $A$ is a many-body quantum system with short-range Hamiltonian $H$, and $\rho=\frac{1}{Z}e^{-\beta H}$ is the thermal equilibrium at temperature $1/\beta$. In that case $\sqrt{\rho}$ is known as the thermofield double (TFD), see [54, 55] for realization proposals on near-term quantum platforms.
The system $A$ undergoes time evolution governed by the Hamiltonian $H$, whereas $\bar{A}$ is left intact. In addition, we perform a measurement on $A$ at $t=\delta t,2\delta t,\dots$. The measurement is described by the Kraus operators $\{K_{m}\}_{m}$ [56], indexed by the possible outcomes $m$, satisfying unitarity $\sum_{m}K_{m}^{\dagger}K_{m}=\mathbf{1}$. Doing this until $t=n\delta t$, we obtain a random sequence of outcomes $\vec{m}=(m_{\delta t},m_{2\delta t},\dots,m_{t})$. In terms of
@@
For each $s=\delta t,2\delta t,\dots,t$, we also introduce the marginal analogues of (2) and (3):
@@
We are ready to define the quantum dynamical entropy:
$S_{\text{CNT}}:=$ $\sum_{\vec{m}}-p_{\vec{m}}\ln p_{\vec{m}}+\sum_{s=\delta t}^{t}\left[\sum_{m}p^{s}_{{m}}\ln p^{s}_{{m}}+J_{s}\right]$
$=$ $S_{\text{cl}}(p)-\sum_{s=\delta t}^{t}\left[S_{\text{cl}}(p^{s})-J_{s}\right]\,.$ (6)
Namely, it is defined as the classical Shannon entropy of the joint outcome distribution $S_{\text{cl}}(p)$, minus a sum of “entropy defects”, $S_{\text{cl}}(p^{s})-J_{s}$, which are the difference between the outcome Shannon entropy and the information gain of a single measurement.
The entropy defect accounts for a key feature of quantum mechanics: measurements inevitably back-react on the system, and can thereby prevent later observations from gaining information about the initial state [19, 16]. For this reason, the Shannon entropy of the measurement outcome alone is not a good generalization of the Kolmolgorov-Sinai entropy to quantum systems, and the entropy defect was introduced in [10] to correct for the measurement backreaction.
Our definition (Planckian bound on quantum dynamical entropy) is a simplification of CNT’s proposal. In that work, the measurements are replaced by a more general notion of “abelian models”, over which a supremum is then taken. Here, we do not take a supremum over the measurement schemes. Indeed, in a many-body system, its choice is nontrivial due to the measurement backreaction effect, as we discuss below.
Choice of observable. To start, we shall restrict to monitoring a single real-valued observable. In particular, we do not perform an extensive number of measurements at each time step. This simplifying restriction is ultimately motivated by the Planckian bound, as the rate can be trivially increased by monitoring multiple weakly correlated regions.
Within this restriction, a natural possibility is a microscopic measurement, which acts on $O(1)$ local degrees of freedom. Note that we assume those to be non-semiclassical, and have a small Hilbert space. We argue that the asymptotic CNT entropy rate cannot be positive for a generic dynamics, due to the significant backreaction caused by microscopic measurements.
In fact, we argue that $J_{t}\to 0$ as $t\to\infty$; from the definition (Planckian bound on quantum dynamical entropy) and the subaddictivity of the Shannon entropy, we have $S_{\text{CNT}}/t\leq\sum_{s\leq t}J_{s}/t$, so $J_{t}\to 0$ would forbid a positive entropy rate. For that, observe from (4), (2), and unitarity, that
@@
The above argument can be adapted to any weak measurements on a few sites, as the dissipative evolution $\mathcal{DL}$ will only preserve the identity operator generically [^1] An exceptional case requiring further discussion is when $K_{m}$ is conserved by the dynamics. In that case, $J_{t}\not\to 0$. However, the outcome Shannon entropy cannot grow linearly, so the dynamical entropy rate still vanishes. Another fine-tuned counter-example is spatial translation [10] where the dynamical entropy rate is nonzero.. Numerically we observe that $J_{t}$ decays exponentially in a non-integrable spin chains, see Fig. 2-(a).
To recapitulate, microscopic measurements are too intrusive to yield a positive entropy rate. We are led to consider measuring a “mesoscopic observable” of the form
@@
where $\left<A\right>:=\mathrm{Tr}[\rho A]$, $Q_{x}$ is a local operator centered around $x$, and $V$ is the volume (number of sites in a lattice model). We shall assume $\left<Q_{x}\right>=0$ upon replacing $Q_{x}\to Q_{x}-\left<Q_{x}\right$>, so that $Q$ describes the fluctuation of an extensive quantity around its ensemble average (hence the term “mesoscopic”). We consider the family of Kraus operators indexed by a continuous outcome
@@
which describe a weak measurement of $Q$ with order-one precision. The definition (Planckian bound on quantum dynamical entropy) can be adapted by replacing sums by integrals; $p_{\vec{m}}$ and $p_{m}^{s}$ become probability densities and the Shannon entropies become differential entropies. The measurement strength can be tuned by rescaling $Q$ itself. The $\sqrt{\delta t}$ factor allows a well-defined continuous monitoring $\delta t\to 0$ limit. Such mesoscopic monitoring causes qualitatively milder disturbance. Numerically, we observe that $J_{t}$ (5) decays only at a time scale $t\sim\sqrt{L}$ in a 1D system of size $L$. Therefore, if we take the thermodynamic limit first, $S_{\text{CNT}}/t$, may have a nonzero $t\to\infty$ limit. Moreover, this limit can be analytically calculated, as we show next.
Entropy rate formula. In the thermodynamic limit and away from thermal criticality, the time-dependent fluctuation of $Q$ becomes Gaussian due to cluster decomposition [57]. As a consequence the dynamical entropy rate can be computed solely in terms of the Keldysh and retarded Green functions in frequency domain [58]. We define them with the following convention:
@@
At thermal equilibrium, they are related by the fluctuation-dissipation theorem (FDT), $\mathrm{Im}G_{R}(\omega)=\tanh(\beta\omega)G_{K}(\omega)/2$; $\mathrm{Re}G_{R}(\omega)$ is given by Kramers-Kronig. In terms of the eigenstate thermalization hypothesis [59, 60, 26], the Green functions encode the same data as the two-point structure factor (yet Gaussianity is not equivalent to the vanishing of higher free cumulants [61, 62]). The Gaussianity enables a simple method to compute the dynamical entropy. We write a free boson model in which a displacement operator has the same $G_{K}$, and perform exact calculation therein. The calculation is lengthy yet elementary; see [57] for details. The result is the following:
@@
where $\tilde{G}(\omega):=G_{K}(\omega)+|G_{R}(\omega)|^{2}$. In particular, the limit of the quantity $J_{t}$ is given by
@@
In Fig. 2-(b), we compare this prediction to numerical data from non-integrable quantum spin chains of moderate sizes, and find nonetheless a good agreement.
Figure 3: Keldysh correlation function $G_{K}$ (as function of $\beta\omega$) that maximizes $\beta s_{\text{CNT}}$ (12). $G_{K}(-\omega)=G_{K}(\omega)$, obtained from numerically maximizing (12) subject to the fluctuation-dissipation theorem. Inset: comparison with $4/|\beta\omega|^{-1}$ (dashed line).
@@
Planckian bound. The entropy rate given by (12) satisfies a Planckian bound. To see this, we make a change of variable $y=\beta\omega$ in its right hand side, which then equals $\beta^{-1}$ times a $\beta$-independent integral, if we view $G_{K}$ and $G_{R}$ as functions of $y$. An explicit but untight upper bound can be obtained by maximizing the integral with respect to $G_{K}$ and $G_{R}$ ignoring FDT. This gives the optimum $|G_{R}|^{2}=0,G_{K}=y/(\sinh y-y)$, and
$\beta s_{\text{CNT}}\leq\int\frac{dy}{4\pi}\left[\ln\frac{\sinh(y)}{\sinh(y)-y}-\frac{y}{\sinh y}\right]\approx 0.57$ (14)
We may obtain a numerically tight bound by solving the FDT-constrained maximization: $\beta s_{\text{CNT}}\leq 0.375\dots$. The optimal $G_{K}$ (see Fig. 3) has a singularity $G_{K}\sim|\beta\omega|^{-\alpha}$ with $\alpha\approx 1$ at $\omega\to 0$.
We observe that in the rate formula (12), the factor $\beta\omega/\sinh(\beta\omega)$ suppresses contributions from $|\omega|\gg 1/\beta$. Thus, at low temperature, the dynamical entropy rate is only sensible to longtime ($\tau\gtrsim\beta$) properties of the monitored dynamics. Therefore, although we derived the exact rate formula for the specific form of the monitoring scheme (10), we expect a large class of monitoring schemes should have nearly identical “coarse-grained” behaviors at time scale $\tau\gtrsim\beta$, and thus obey also the Planckian bound. Furthermore, recall also that microscopic monitoring generically yields vanishing dynamical entropy rate. Motivated by these considerations, we conjecture that the dynamical entropy rate obeys a universal Planckian bound $\beta s_{\text{CNT}}\leq c$, with a universal constant $c$, and with the exception of thermal critical points, for any single-operator monitoring.
Intuitively, the Planckian bound can be explained as follows. The information on $\bar{A}$ accessible from $A$ is encoded in correlation functions across the TFD, of the form $\mathrm{Tr}[O\sqrt{\rho}Q(t)\sqrt{\rho}]$. Their Fourier transform is known to decay as $e^{-\beta|\omega|/2}$ or faster [32, 43]. This origin of the Planckian bound is reminiscent of that on the OTOC growth rate (Lyapunov exponent) [37], whose derivation often requires “regularizing” the OTOC as a correlation across the TFD, of form $\mathrm{Tr}[[A(t),B]\sqrt{\rho}[A(t),B]\sqrt{\rho}]$ [37, 63, 64] (see however [65, 66, 67] for discussions on the necessity of the regularization). In our setup, the TFD occurs naturally in order to encode the information about the initial thermal state.
Purification. Using the same monitoring setup of Fig. 1, one may define purification (also known as conditional entropy), as the decrease of the von Neumann entropy of $\bar{A}$ induced by the measurements, averaged over the outcomes (weighed by Born’s rule):
$J:=\sum_{\vec{m}}p_{\vec{m}}\mathrm{Tr}[\rho_{\vec{m}}\ln\rho_{\vec{m}}-\rho\ln\rho]\,.$ (15)
$J$ quantifies the amount of initial condition revealed by the observer [68, 69]. It is also a probe of measurement-induced phases  [70, 71, 72, 73, 74]. We find that the purification rate of $J/t$, under mesoscopic monitoring (9)-(10), is also given by an exact formula,
@@
$\mathcal{J}=\lim_{\begin{subarray}{c}t\to\infty\\ \delta t\to 0\end{subarray}}\lim_{V\to\infty}\frac{J}{t}=\int\frac{d\omega}{4\pi}\frac{\beta\omega}{\sinh\beta\omega}\frac{G_{K}}{1+G_{K}+|G_{R}|^{2}},$ (16)
@@
derived under the same assumption and with the same method [57]. We checked (16) with exact numerical simulations in small systems and find an excellent agreement, see Fig. 2-(d). Eq. (16) also satisfies a Planckian bound: since $G_{K}\geq 0$ and $|G_{R}|^{2}\geq 0$ are positive, $G_{K}/(1+G_{K}+|G_{R}|^{2})\leq 1$, and $\beta\mathcal{J}\leq\pi/8$, for a mesoscopic observable. For microscopic monitoring, we observe numerically that $J/t\to 0$ as well; in fact, $J\sim\ln t$ in our example, see Fig. 2-(c). Therefore we conjecture that single-measurement purification also satisfies a general Planckian bound. In summary, we find that purification has similar properties as dynamical entropy, while being slightly simpler and more accessible numerically.
Conclusion. We showed that the quantum dynamical entropy (and purification, same below) can have a well-defined linear growth in generic quantum many-body systems, and proposed a Planckian bound for its growth rate. Our results concern the time regime where $t$ is large compared to the inverse coupling such as the Thouless time, or the time scale $t\sim e^{cV}$ where purification becomes model-independent [75, 76]. In this regime, the dynamical entropy rate with mesoscopic monitoring depends only on linear response quantities, and is not a direct probe of integrability, although integrability often leads to distinct linear response properties and influences dynamical entropy indirectly. A natural question for future work would be to understand contributions from nonlinearity at some system-size dependent time scale, and make connection with emergent hydrodynamics [7, 77, 78, 79, 80, 81, 81, 82].
We have focused on systems away from semiclassical or large $N$ limit, yet our results do apply to large $N$ models like SYK , where a comparison with OTOCs would be useful. More generally, we would like to identify low-temperature states with large dynamical entropy. A final challenge lift the “single-operator” restriction in a sensible way [83], and propose a quantum many-body extension of the Pesin relation [17, 84, 85, 86].

### v4

@@ Abstract
We introduce a simplified version of Connes-Narnhofer-Thirring’s quantum dynamical entropy for quantum systems. It quantifies the amount of information gained about the initial condition from continuously monitoring an observable. A nonzero entropy growth rate can be obtained by monitoring the thermal fluctuation of an extensive observable in a generic many-body system, away from classical or large $N$ limits. We explicitly compute the entropy rate in the thermodynamic and long-time limit, in terms of the two-point correlation functions. We conjecture a universal Planckian bound for the entropy rate. Related results on the purification rate are also obtained.
@@
Defining a dynamical entropy for quantum systems has been an outstanding question since KS [8]. A few proposals have been put forward [9, 10, 11, 12, 13, 14, 15, 16, 17], and concrete calculations have been performed for simple systems such as free quantum gases [18, 14, 19]. However, it remains unknown whether quantum dynamical entropy can characterize chaos and complexity in many-body quantum systems (see [20] for a recent survey). This question is timely as many-body quantum chaos has been an intensively pursued topic [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. One interesting approach is the out-of-time order correlation (OTOC) [37],
$\text{OTOC}=\mathrm{Tr}[\sqrt{\rho}[P,Q(t)]^{\dagger}\sqrt{\rho}[P,Q(t)]]$ (1)
where $\rho=e^{-\beta H}/\mathrm{Tr}[e^{-\beta H}]$ is the thermal state at temperature $1/\beta$ of some quantum system with Hamiltonian $H$, $P$ and $Q$ are operators, and $Q(t)=e^{iHt}Qe^{-iHt}$. OTOCs generalize the classical butterfly effect, that is, the sensitivity of an observable at $t$ to initial perturbations, to quantum systems. Remarkably, its growth rate (Lyapunov exponent) obeys a Planckian bound at low temperatures [37],
$\text{OTOC}\sim e^{\lambda_{L}t}\implies\lambda_{L}\leq 2\pi T,T=1/\beta.$ (2)
The Planckian bound showcases a universal limit on dynamical chaos due to quantum effects, and is famously saturated by strongly interacting holographic models [38, 39, 40, 41, 42]; see also [43, 44, 45, 46, 47, 48] for other “Planckian bounds”. However, OTOCs do not have a well-defined exponential growth in generic many-body quantum systems with local interactions [22, 49, 50], that is, away from semiclassical or large $N$ limits. Meanwhile, quantum chaos diagnosis inspired by random matrix theory [51, 52, 53] probes time scales that diverge in the thermodynamic limit.
In this Letter, we propose a quantum chaos probe that remedies the above issues. It is a simplified version of the Connes-Narnhofer-Thirring (CNT) quantum dynamical entropy [10]. Like the OTOC, it is defined for a thermal state $\rho$ and an operator $Q$. Unlike OTOC, it involves repeatedly measuring (monitoring [19, 54, 55]) $Q$ from 0 to $t$, and is defined as follows,
$S_{\text{CNT}}:=S_{\text{cl}}(p)-\sum_{s\leq t}\left[S_{\text{cl}}(p^{s})-J_{s}\right]\,$, (3)
where $S_{\text{cl}}(p)$ is the classical Shannon entropy of the outcome distribution, $S_{\text{cl}}(p^{s})$ that of the marginal distribution at time $s$, and $J_{s}$ the amount of information gained by the measurement at $s$; see below and in particular (8)-(10) for more details. The CNT entropy may be more accessible in experiments than recently studied measurement-induced quantities [56, 57], as it involves a milder post-selection problem.
Remarkably, for suitable $Q$ and measurement scheme, $S_{\text{CNT}}$ has a well-defined linear growth, in generic quantum many-body systems, and at a time scale independent of the system size (and only determined by the local energy scale). Moreover, we conjecture that the growth rate satisfies a Planckian bound:
@@
where $V$ is the volume, and $c$ is a universal constant. We support the conjecture (4) with two main pieces of evidence. First, its growth rate vanishes if $Q$ is a local observable. Second, if $Q$ is a rescaled extensive quantity [“mesoscopic observable”, (13) below], the entropy growth rate is given, in a large non-critical many-body systems and for a Gaussian measurement scheme [see (14) below], by an exact formula [see (16) below], which satisfies a Planckian bound [see (18) below]. The exact rate formula is obtained thanks to a general emergent Gaussianity property of interacting systems, and is our main technical contribution. We shall then argue that the Planck bound applies more generally.
Below, we shall define our setup in more detail, see Fig. 1, go through the arguments supporting the Planckian bound, and discuss analogous results on purification.
@@
Setup. Consider a many-body quantum system $A$ with short-range Hamiltonian $H$, initialized in a thermal state $\rho=e^{-\beta H}/\mathrm{Tr}[e^{-\beta H}]$ at temperature $1/\beta$. We purify $\rho$ as the reduced density matrix of the pure state $\sqrt{\rho}$ on $A\bar{A}$, where $\bar{A}$ is an identical copy of $A$. The state $\sqrt{\rho}$ is known as the thermofield double (TFD), see [58, 59] for realization proposals on near-term quantum platforms.
The system $A$ undergoes time evolution under $H$, whereas $\bar{A}$ is left intact. In addition, we monitor an operator $Q$ acting on $A$ at $t=\delta t,2\delta t,\dots$. The measurement is described by a family of Kraus operators $\{K_{m}\}_{m}$ depending on $Q$, indexed by possible measurement outcomes $m$, and satisfying unitary $\sum_{m}K_{m}^{\dagger}K_{m}=\mathbf{1}$ [60]. For example, we may take
$K_{m}=(\mathbf{1}+mQ)/\sqrt{2(\mathbf{1}+m^{2}Q^{2})},m=\pm 1,$
or (14) below. Doing this until $t=n\delta t$, we obtain a random sequence of outcomes $\vec{m}=(m_{\delta t},m_{2\delta t},\dots,m_{t})$. In terms of
@@
In terms of these, purification is defined as
$J:=\sum_{\vec{m}}p_{\vec{m}}\mathrm{Tr}[\rho_{\vec{m}}\ln\rho_{\vec{m}}-\rho\ln\rho]\,$, (7)
Namely, it equals how much the continuous monitoring reduces the entanglement between $A$ and $\bar{A}$. $J$ quantifies the amount of information (on the initial condition) revealed by the measurements [61, 62], and has been used as a probe of measurement-induced phases [63, 64, 65, 66, 67].
For the CNT entropy, we introduce, for each $s=\delta t,2\delta t,\dots,t$, the marginal analogues of (5), (6), and (7),
@@
Then the CNT entropy’s definition (3) is completed by the recalling the definition of Shannon entropies,
$S_{\text{cl}}(p):=\sum_{\vec{m}}-p_{\vec{m}}\ln p_{\vec{m}},S_{\text{cl}}(p^{s}):=\sum_{{m}}-p^{s}_{{m}}\ln p^{s}_{{m}}.$ (10)
The CNT entropy is in principle accessible experimentally. $S_{\text{cl}}(p)$ is the Shannon entropy of a random time sequence (sampled by experimental runs), and can be estimated by classical statistical methods [68]. $S_{\text{cl}}(p^{s})$ concerns 1D distributions, so is simpler to estimate. In $J_{s}$, the entanglement entropies can be estimated by modern tomography techniques [69]. Crucially, there is not a hard post-selection problem: Since we condition on a single outcome $m_{s}$, we do not need $O(e^{t})$ number of experimental runs to achieve one outcome many times in order to perform tomography. Purification is more challenging, as it involves an $O(e^{t})$ hard post-selection [56, 57, 70].
Our choice to monitor a single observable $Q$ is motivated by the comparison with OTOC, and ultimately by the Planckian bound, as the entropy can be trivially increased by monitoring multiple uncorrelated regions. This makes our definition simpler than CNT’s proposal, which involves maximizing over measurement schemes.
Nevertheless, we kept the key idea of CNT, which is the term $S_{\text{cl}}(p^{s})-J_{s}$ in (10), known as the “entropy defect”. It accounts for a feature of quantum mechanics: Measurements backreact on the system, and can thus prevent later observations from gaining information about the initial state [19, 16]. For this reason, $S_{\text{cl}}(p)$ alone is not a good definition of quantum dynamical entropy. Also, the growth of $S_{\text{CNT}}$ depends qualitatively on the nature of the operator $Q$, as we discuss next.
Local operators. We first consider the case where $Q$ is a local operator acting on $O(1)$ local degrees of freedom. We assume those to be non-semiclassical, and have a small Hilbert space. We argue that the asymptotic CNT entropy rate cannot be positive for a generic dynamics, due to the significant backreaction (dephasing).
In fact, we argue that $J_{t}\to 0$ as $t\to\infty$; from the definition (3) and the subaddictivity of the Shannon entropy, we have $S_{\text{CNT}}/t\leq\sum_{s\leq t}J_{s}/t$, so $J_{t}\to 0$ would forbid a positive entropy rate. For that, observe from (8), (5), and unitarity, that
@@
The above argument can be adapted to any weak measurements on a few sites, as the dissipative evolution $\mathcal{DL}$ will only preserve the identity operator generically [^1] An exceptional case is when $Q$ is conserved by the dynamics. In that case, $J_{t}\not\to 0$. However, since $Q(t)=Q$, the outcome Shannon entropy cannot grow linearly, so the dynamical entropy rate still vanishes. Another fine-tuned counter-example is spatial translation [10] where the dynamical entropy rate is nonzero.. Numerically we observe that $J_{t}$ decays exponentially in a non-integrable spin chain, see Fig. 2-(a).
Mesoscopic operators. As local operators cannot yield a positive entropy rate, we are led to consider measuring a “mesoscopic observable” of the form
@@
where $\left<A\right>:=\mathrm{Tr}[\rho A]$, $Q_{x}$ is a local operator centered around $x$, and $V$ is the volume (number of sites in a lattice model). We may always assume $\left<Q_{x}\right>=0$ upon replacing $Q_{x}\to Q_{x}-\left<Q_{x}\right$>, so that $Q$ describes the fluctuation of an extensive quantity around its average (hence the term “mesoscopic”). We consider the family of Kraus operators indexed by a continuous outcome
@@
which describe a weak measurement of $Q$ with order-one precision. The definition (10) can be adapted by replacing sums by integrals; $p_{\vec{m}}$ and $p_{m}^{s}$ become probability densities and the Shannon entropies become differential entropies. The measurement strength can be tuned by rescaling $Q$ itself. The $\sqrt{\delta t}$ factor allows a well-defined continuous monitoring $\delta t\to 0$ limit. Such mesoscopic monitoring has a qualitatively milder backreaction, and can lead to a nonzero entropy growth rate.
Rate formula. In fact, the aforementioned rate can be analytically calculated, even in interacting systems. Indeed, in the thermodynamic limit and away from thermal criticality, the time-dependent fluctuation of $Q$ becomes effectively Gaussian due to cluster decomposition (see End Matter). Thus, the dynamical entropy rate can be computed solely in terms of the Keldysh and retarded Green functions in frequency domain [72]. We define them with the following convention:
@@
At thermal equilibrium, they are related by the fluctuation-dissipation theorem (FDT), $\mathrm{Im}G_{R}(\omega)=\tanh(\beta\omega)G_{K}(\omega)/2$; $\mathrm{Re}G_{R}(\omega)$ is given by Kramers-Kronig. In terms of the eigenstate thermalization hypothesis [73, 74, 26], the Green functions encode the same data as the two-point structure factor (yet Gaussianity is not equivalent to the vanishing of higher free cumulants [75, 76]). The emergent Gaussianity enables a simple method to compute the dynamical entropy. We write a free boson model in which a displacement operator has the same $G_{K}$, perform exact calculation therein. The calculation is lengthy yet elementary; see [71] for details. The result is the following:
@@
where $\tilde{G}(\omega):=G_{K}(\omega)+|G_{R}(\omega)|^{2}$. In particular, the limit of the quantity $J_{t}$ is given by [77],
@@
We emphasize that (16) and (17) are exact for interacting systems, not just free bosons, thanks to emergent Gaussianity. In Fig. 2-(b), we compare this prediction to numerics in a non-integrable quantum spin chain for moderate length $L$. Although the numerical value deviates from the prediction as $t$ increases for any fixed $L,\delta t$, we observe a single-variable scaling $J_{t}(t,{L},\delta t)=g(t/\sqrt{L},\delta t)$ where $g(0,\delta t)$ tends to the prediction as $\delta t\to 0$. So, if we send $L\to\infty$ and then $t\to\infty,\delta t\to 0$, as in (17), we find a nice agreement between theory and numerics.
@@
Planckian bound. The entropy rate given by (16) satisfies a Planckian bound. To see this, we make a change of variable $y=\beta\omega$ in its right hand side, which then equals $\beta^{-1}$ times a $\beta$-independent integral, if we view $G_{K}$ and $G_{R}$ as functions of $y$. An explicit upper bound is obtained by maximizing the integral with respect to $G_{K}$ and $G_{R}$ ignoring FDT. This gives the optimum $|G_{R}|^{2}=0,G_{K}=y/(\sinh y-y)$, and
$\beta s_{\text{CNT}}\leq\int\frac{dy}{4\pi}\left[\ln\frac{\sinh(y)}{\sinh(y)-y}-\frac{y}{\sinh y}\right]\approx 0.57.$ (18)
See End Matter for a tighter bound taken into account FDT, and Fig. 3 for the optical $G_{K}$.
We conjecture that the Planckian bound applies to any many-body system away from thermal criticality, and for a general class of operators and measurement schemes. We just showed it for extensive operators rescaled by $1/\sqrt{V}$ and the Gaussian scheme (14). We showed also that more local operators in general yield a vanishing dynamical entropy rate. Changing the rescaling factor $1/\sqrt{V}$ to another power would amount to making $G_{K},G_{R}$ very big or small in (16), and that cannot break bound (18): The optimum has $G_{K},G_{R}\sim O(1)$. Finally, we cannot prove, but we surmise the Gaussian scheme (14) to be a renormalization group (RG) fixed point of a large class of schemes under temporal coarse-graining. So, all these schemes should satisfy a universal Planckian bound in the most nontrivial low-$T$ limit [Planckian bounds like (2) and (18) are usually far from tight at high-$T$ since the growth rate is limited by the local energy scale and does not diverge with $T$.] In light of these arguments, the Planckian bound (4) is a rather plausible conjecture.
We remark that in the rate formula (16), the factor $\beta\omega/\sinh(\beta\omega)$ is exponentially small for $|\omega|\gg 1/\beta$. So, at low temperature, only slow fluctuations beyond the “Planckian” timescale $\tau\gtrsim\beta$ contribute to the dynamical entropy. Indeed, the information on $\bar{A}$ accessible from $A$ is encoded in correlation functions across the TFD, of the form $\mathrm{Tr}[Q\sqrt{\rho}Q(t)\sqrt{\rho}]$, which are known to decay as $e^{-\beta|\omega|/2}$ or faster at large $\omega$ [32, 43]. In this regard, our Planckian bound resembles that of OTOC, whose derivations [37, 78, 79] often rely on the TFD in (1); see however [80, 81, 82].
Purification. Under the same assumptions and using the same methods as (16), we may obtain an exact formula for the purification rate [71],
@@
$\mathcal{J}:=\lim_{\begin{subarray}{c}t\to\infty\\ \delta t\to 0\end{subarray}}\lim_{V\to\infty}\frac{J}{t}=\int\frac{d\omega}{4\pi}\frac{\beta\omega}{\sinh\beta\omega}\frac{G_{K}}{1+G_{K}+|G_{R}|^{2}}.$ (19)
@@
This results complements studies of purification at time $t\sim e^{cV}$ [83, 84]. We checked (19) with exact numerical simulations in small systems and find an excellent agreement, see Fig. 2-(d). Eq. (19) also satisfies a Planckian bound: Since $G_{K},|G_{R}|^{2}\geq 0$, we have $G_{K}/(1+G_{K}+|G_{R}|^{2})\leq 1$, and $\beta\mathcal{J}\leq\pi/8$. For local observables, we observe numerically that $J/t\to 0$ as well; in fact, $J\sim\ln t$ in our example, see Fig. 2-(c). So we conjecture that purification also satisfies a general Planckian bound. In summary, purification has similar properties as dynamical entropy, while being simpler to measure numerically (but not experimentally).
Conclusion. We proposed probes of low-energy quantum chaos which have promising behaviors for generic quantum many-body systems away from semiclassical and large $N$ limits, and worth further theoretical or experiment study.
Under quite general conditions, these probes depend only on linear response functions, and do not distinguish directly integrable and non-integrable systems (unlike random-matrix inspired probes [53, 52]). Indeed, one can reproduce any linear response function in a free system. Yet, while not probed directly, integrability often leads to distinct linear response properties and influences dynamical entropy indirectly. It will be useful to study this point in detail.
We close with a few other directions for future work. We would like to understand corrections to the emergent Gaussianity, and explore connection with nonlinear hydrodynamic tails [7, 85, 86, 87, 88, 89, 89, 90]. We should also identify low-temperature states with large dynamical entropy, and compare quantitatively with OTOC’s in large $N$ systems. Finally, we wish to extend our work to monitoring multiple operators monitoring [54], and propose a quantum many-body extension of the Pesin relation [17, 91, 92, 93].